The base of an isosceles triangle 10 feet long and the base angles are decreasing at a rate of 2 degress per second. Find the rate of change of the area when the base angles are 45 degrees?


1
Expert's answer
2021-01-19T18:08:15-0500


Here in the figure BC=10ft. And we assume "\angle B=\angle C =\theta" (cause isosceles triangle).


Also BD=DC=5ft (cause AD is the perpendicular bisector of BC).


Now, in "\triangle ABD" , using trigonometry:-



"\tan(\angle ABD)=\frac{AD}{BD}\\~\\
\Rightarrow\tan\theta=\frac{AD}{5}\\~\\
\Rightarrow AD=5\tan\theta"

So, now we calculate the area of the "\triangle ABC" :-



"\triangle=\frac{1}{2}\cdot5\tan\theta\cdot10\\~\\
\Rightarrow\triangle=25\tan\theta\\"

Now, as we're dealing with rate of change of different stuff we differentiate both sides wrt "t" :-



"\frac{d\triangle}{dt}=25\sec^2\theta\frac{d\theta}{dt}\\~\\
\Rightarrow \frac{d\triangle}{dt}=25\cdot(\sqrt{2})^2\cdot2\\~\\
\Rightarrow \frac{d\triangle}{dt}=100"

We, put "\theta=45^\degree" , and use the rate of change of "\theta" given in the question.


So, rate of change of area is "100" sq ft per second.